Finch Chicken Corporation processes and packages chicken for grocery stores. It purchases chickens from farmers and processes them into two different products: chicken drumsticks and chicken steak. From a standard batch of 15,000 pounds of raw chicken that costs $7,900, the company produces two parts: 3,000 pounds of drumsticks and 6,000 pounds of breast for a processing cost of $6,491. The chicken breast is further processed into 5,200 pounds of steak for a processing cost of $2,000. The market price of drumsticks per pound is $1.55 and the market price per pound of chicken steak is $4.00. If Finch decided to sell chicken breast instead of chicken steak, the price per pound would be $2.30.

Required

  1. a-1. Allocate the joint cost to the joint products, drumsticks and breasts, using weight as the allocation base.

  2. a-2. Calculate the gross margin for each product.

  3. a-3. If the drumsticks are producing a loss, should that product line be eliminated?

  4. b-1. Reallocate the joint cost to the joint products, drumsticks and breasts, using relative market values as the allocation base.

  5. b-2. Calculate the gross margin for each product.

  6. c-1. Should Martin further process chicken breasts into chicken steak? (Use the assumption made in requirement b-1).

  7. c-2. How would the profit be affected by your answer in c-1?
Answer #1

Solution a1:

Total joint cost = $7,900 + $6,491 = $14,391

Joint cost allocated to drum sticks = $14,391 * 3000/9000 = $4,797

Joint cost allocated to breast = $14,391 * 6000 / 9000 = $9,594

Solution a2:

Computation of Gross Margin
Particulars Drumsticks Breast
Sales $4,650.00 $13,800.00
Allocated costs $4,797.00 $9,594.00
Gross margin -$147.00 $4,206.00

Solution a3:

If drumstick are producing loss, product line should not be eliminated as another product is producing profit.

Solution b1:

Allocation of joint costing using relative market value:

Joint cost allocated to drum sticks = $14,391 * 4650/18450 = $3,627

Joint cost allocated to breast = $14,391 * 13800 / 18450 = $10,764

Solution b2:

Computation of Gross Margin
Particulars Drumsticks Breast
Sales $4,650.00 $13,800.00
Allocated costs $3,627.00 $10,764.00
Gross margin $1,023.00 $3,036.00

Solution c1:

Computation of incremental profit on further processing of chicken breast
Particulars Amount
Sale value after further processing (5200*$4) $20,800.00
Sale value at split off point $13,800.00
Incremental revenue $7,000.00
Further processing cost $2,000.00
Incremental profit on further processing $5,000.00

Yes, martin should further process chicken breasts into chicken steak

Solution c2:

Profit will increase by $5,000